OpenStudy (nuccioreggie):

Naomi works in a music shop and earns $42 per day. She earns an extra $2 for each CD she sells. If Naomi wants to earn at least $80 per day, which inequality shows the minimum number of CDs, n, that she should sell? 42 + 2n ≤ 80, so n ≤ 19 42 + 2n ≥ 80, so n ≥ 19 42 + 2n ≥ 80, so n ≥ 36 42 + 2n ≤ 80, so n ≤ 36

OpenStudy (nuccioreggie):

@dan815

OpenStudy (nuccioreggie):

@horsegirl325

OpenStudy (nuccioreggie):

@confluxepic

OpenStudy (nuccioreggie):

@johnweldon1993

OpenStudy (anonymous):

the third one

OpenStudy (nuccioreggie):

@surjithayer

OpenStudy (johnweldon1993):

Well, she wants to earn AT LEAST 80...so we wont settle for less so we need to have a \(\large \ge\) sign in our answer...so that gets rid of A and D right?

OpenStudy (anonymous):

at least 80 means 80 or more so \[42+2n \ge 80 \] solve for n

OpenStudy (nuccioreggie):

Yes thats y i thought it was C

OpenStudy (anonymous):

no \[2n \ge 80-42,2n \ge 38,n \ge19\]

OpenStudy (nuccioreggie):

Oooooooooo ok it B

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

guys it is agianst openstuudy code of conduct to give out answers u must let them learn it